OpenStudy (anonymous):

By what number should you multiply the first equation to solve using elimination? –4x – 2y = –24 –8x + 3y = –34

OpenStudy (anonymous):

2

OpenStudy (anonymous):

(-4x-2y=-24) -8x-4y=-48

OpenStudy (anonymous):

Thank you!